# Leveraging MTD to Mitigate Poisoning Attacks in Decentralized FL with Non-IID Data

### v2

@@ Abstract
Decentralized Federated Learning (DFL), a paradigm for managing big data in a privacy-preserving and distributed manner, is still vulnerable to poisoning attacks where malicious clients tamper data or models. Current defense methods often assume Independently and Identically Distributed (IID) data across participants, which is unrealistic in real-world applications. In more realistic non-IID contexts, existing defensive strategies face challenges when distinguishing between models that have been compromised and those that have been trained on heterogeneous data distributions (non-IID), leading to diminished efficacy. In response, this paper proposes a framework that employs the Moving Target Defense (MTD) approach to bolster the robustness of DFL models. By continuously modifying the attack surface of the DFL system, the framework aims to mitigate poisoning attacks effectively. The proposed solution includes both proactive and reactive modes, utilizing a reputation system that combines metrics of model similarity and loss, alongside various defensive techniques. Comprehensive experimental evaluations indicate that the MTD-based mechanism significantly mitigates a range of poisoning attack types across multiple datasets with different federation topologies.
@@ I Introduction
This work proposes an MTD-based approach to dynamically modify the attack surface of DFL systems and improve their resilience and security. Therefore, the main contributions of this paper are: (1) Designing an MTD-based approach to mitigate poisoning attacks in DFL. This approach utilizes a dual-reputation score system based on model similarity and loss, adopts proactive and reactive modes, and incorporates dynamic topology and aggregation strategies. (2) Employing the DBSCAN clustering ‘algorithm to adjust the triggering threshold for MTD dynamically. (3) Implementing a prototype of the proposed MTD method on a real DFL platform [11]. (4) Conducting extensive experiments to compare the proposed algorithm with state-of-the-art methods, including scenarios with fully connected, ring, and star topologies, as well as untargeted poisoning and backdoor attacks on MNIST, FashionMNIST, and CIFAR-10 datasets. The results demonstrate that the proposed MTD-based algorithm not only effectively mitigates poisoning attacks, outperforming existing methods, but also improves model performance on non-IID data.
@@ II Background and Related Work
[14] 2020 ML Shuffling Model Weights Proactive Adversarial Attack
@@
TABLE I provides an overview of the existing literature on applying MTD strategies in the context of Artificial Intelligence (AI) Security to address diverse cyber threats. In the traditional AI field, MTD has been employed to mitigate adversarial samples and poisoning attacks. [12] proposes an MTD approach to enhance Deep Neural Network (DNN) robustness by dynamically changing model parameters during inference. This makes it harder for attackers to craft compelling adversarial examples, improving resilience against white and black-box attacks. Morphence [13] is an MTD strategy generating diverse student models from a base model with random weight perturbations. Randomly selecting a student model for each inference reduces the adversarial attack success rate while maintaining accuracy. EI-MTD [14] is a dynamic defense mechanism for edge computing. It uses a three-stage framework to address challenges like limited resources and transferability, aiming to prevent adversaries from finding substitute models and reducing attack transferability. MT-MTD [15] presents an MTD framework protecting DNNs against backdoor attacks in edge computing. It divides training data, has an attacker train each dimension, and the defender randomly selects dimensions for consensus, increasing attack cost and maintaining accuracy. MTDeep [16] enhances neural network robustness using MTD by dynamically changing model architecture, making it harder for attackers to craft adversarial attacks.
However, there is limited research on utilizing MTD methodology to address the security challenges of FL systems. [17] enhances FL security with MTD by introducing an augmented dual-shuffle mechanism. This dynamically changes the attack surface by shuffling client selection and model aggregation, improving system security against various poisoning attacks. [9] introduces a proactive MTD mechanism that continuously reconfigures the network topology to enhance DFL security. Voyager [10] employs a reactive MTD-based aggregation protocol to bolster the resilience of DFL against poisoning attacks. This is achieved by dynamically manipulating the connectivity of the network topology. However, its efficacy remains to be validated in non-IID scenarios.
@@ IV-B MTD Module
An MTD module can be characterized by a triplet $<C,M,T>$, where $C$ signifies the configuration space within which the MTD module is capable of moving, $M$ indicates the strategy employed for the action, and $T$ represents the timing function that determines the intervals at which actions are executed [20]. In essence, the design of an MTD module necessitates addressing three fundamental questions: (1) What to move; (2) How to move; and (3) When to move [21].
What to Move. The threat model analysis delineated in Section III-B reveals that the attackers’ knowledge and capabilities are restricted to their local area. Thus, the attack surface for benign nodes mainly comes from the model shared by the malicious nodes and the aggregation functions. This paper posits that the configurations of neighboring connections in benign nodes and aggregation functions can be dynamically moved to mitigate the risks posed by poisoning attacks. By modifying the selection of neighboring nodes, benign nodes can effectively diminish the influence exerted by malicious nodes, thereby altering the network topology of the federation. Additionally, by varying the aggregation function, the complexity of executing a successful attack targeting a specific aggregation method can be significantly increased, thereby enhancing the system’s resilience against such threats.
@@
How to Move. ”How to move” refers to the principles and strategies used to alter system configuration elements to increase unpredictability and confuse attackers. This work develops two MTD strategies: the dynamic topology strategy and the dynamic aggregation algorithm strategy.
@@ IV-B2 Dynamic Aggregation Algorithm Strategy
When to Move. In the MTD module, ”when to move” pertains to selecting the optimal moment to execute the strategies, transitioning from the existing state to an alternative configuration. This work proposes two modes of MTD: proactive and reactive. In the proactive mode, the benign node modifies its connected neighbors and the aggregation algorithm in each round. Conversely, in the reactive mode, the benign node requires a triggering event before initiating an appropriate response.
In conventional reactive MTD systems, the trigger of a strategy is primarily contingent upon a pre-established threshold [10]. However, this pre-established threshold necessitates a priori knowledge, which is often difficult to achieve in practice. Therefore, this paper proposes a dynamic threshold determination algorithm to automate acquiring the threshold for triggering MTD strategies. As shown in Algorithm 3, this work utilizes the Density-Based Spatial Clustering of Applications with Noise (DBSCAN) [22] algorithm to dynamically find the model similarity threshold, loss threshold, and the MTD trigger.
@@ V-A Deployment Configuration
Dataset MNIST, FashionMNIST, CIFAR-10
@@
  • MNIST [19] is a widely-used handwritten digits (0-9) dataset for image recognition, consisting of 60,000 training images and 10,000 testing images, each in grayscale and sized 28x28 pixels. A Multilayer Perceptron (MLP) model is used, consisting of two fully connected hidden layers with 256x128. The MLP is trained for 3 epochs per federated round, optimized using Adam with a learning rate of $1\mathrm{e}{-3}$, and cross-entropy loss is applied.
  • FashionMNIST [23] is a dataset created as a more challenging alternative to MNIST for benchmarks. It contains 60,000 training and 10,000 testing 28x28 grayscale images across 10 fashion categories (e.g., trousers and coats). A small CNN with two convolutional layers (32 and 64 filters) is employed, followed by two fully connected layers. The model is trained on FashionMNIST for 3 epochs per round using Adam optimizer with a learning rate of $1\mathrm{e}{-3}$.
  • CIFAR-10 [24] is a benchmark dataset for ML, consisting of 50,000 training and 10,000 testing images (32x32 pixels) across 10 classes (e.g., airplane, automobile, bird). The colorful images and diverse classes make it more challenging than MNIST and FashionMNIST. The SimpleMobileNet [25] model is utilized. It starts with a convolutional layer (32 filters, 3x3 kernel), followed by max pooling. Then, five depthwise separable convolutional layers are used, followed by adaptive average pooling and a dense layer. The model is trained using Adam optimizer (learning rate $1\mathrm{e}{-3}$) for 5 epochs per round.
@@
Federation Setups. The experiments were performed using a federation with 10 nodes with 10 rounds. The initial network topology varied among fully connected ring and star structures, as depicted in Figure 3. The network topology could change dynamically during training if the topology MTD was enabled.
Figure 3: Initial Network Topology for DFL
@@ V-B Experiments
alpha Agg. Fully Ring Star Fully Ring Star Fully Ring Star
@@
Figure 4: Average F1-Score Results for Untargeted Label Flipping
@@
Given that the attack strategy employed in this experiments is a random weight attack, the dynamic aggregation algorithm fails to demonstrate its advantage. To summarize, the defense strategies based on MTD prove to be an effective means of mitigating model poisoning attacks within non-IID environments.
@@ VI Discussion
Extensive experiments have demonstrated that the MTD-based approach proposed in this paper can effectively mitigate poisoning attacks in non-IID environments, particularly for untargeted attacks. Therefore, the algorithm presented herein emerges as a reliable defense mechanism. Nevertheless, despite this approach demonstrating superiority over existing state-of-the-art algorithms, there is potential for enhancement in addressing targeted attacks, such as backdoor attacks.
To investigate the reasons behind the sub-optimal performance of the MTD-based approach in backdoor attacks, this work analyzes the distribution of reputation scores generated by benign nodes. It is conducted under conditions where the PNR is set at 10% and within a fully connected topology. The comparison is performed across three datasets, focusing on model poisoning and backdoor attack scenarios.
@@
It is essential to highlight that the proposed solution enhances the model performance of DFL in non-IID environments, especially when using the dynamic topology strategy. The average F1-Score achieved by the proposed approach reaches 0.95 for the MNIST dataset, 0.83 for FashionMNIST, and 0.70 for Cifar10 when using the reactive dynamic topology strategy, as illustrated in Figure4 and 5. These results indicate that the algorithm not only mitigates the impacts of poisoning attacks but also enhances model performance in non-IID settings. The effectiveness of the dynamic topology lies in its tendency to facilitate connections and aggregations among nodes that exhibit high similarity, thereby diminishing the impact of non-IID data.
@@ VII Conclusion and Future Work
This work presents an MTD-based approach to mitigate poisoning attacks in DFL, particularly in non-IID environments. The approach employs a novel dual-reputation score mechanism that evaluates both model similarity and loss. It incorporates proactive and reactive MTD modes coupled with dynamic topology and aggregation algorithm strategies. By utilizing the DBSCAN clustering algorithm, the proposed MTD strategy overcomes the limitations of traditional defense mechanisms that often rely on fixed thresholds. Instead, it dynamically adjusts thresholds and triggers based on the current state of the federation. Furthermore, the dynamic topology strategy promotes connections among nodes that are closely aligned with the local model, enhancing performance in non-IID scenarios. The effectiveness of this MTD-based approach is demonstrated through experiments on MNIST, FashionMNIST, and CIFAR-10 datasets, employing various initial topologies (fully connected, ring, and star) and comparing against label flipping, model poisoning, and backdoor attacks. The results show that the proposed approach effectively mitigates a range of untargeted attacks and outperforms state-of-the-art aggregation algorithms in defending against backdoor attacks. Consequently, this MTD-based approach represents a robust and reliable security framework for non-IID data.

### v3

@@ Abstract
Decentralized Federated Learning (DFL), a paradigm for managing big data in a privacy-preserving and distributed manner, is vulnerable to poisoning attacks where malicious clients tamper with data or models. Current defense methods often assume Independently and Identically Distributed (IID) data across participants, which is unrealistic in real-world applications. In more realistic non-IID contexts, existing defensive strategies face challenges when distinguishing between models that have been compromised and those that have been trained on heterogeneous data distributions (non-IID), leading to diminished efficacy. In response, this paper proposes a framework that employs the Moving Target Defense (MTD) approach to bolster the robustness of DFL models. By continuously modifying the attack surface of the DFL system, the framework aims to mitigate poisoning attacks effectively. The proposed solution includes both proactive and reactive modes, utilizing a reputation system that combines metrics of model similarity and loss, alongside various defensive techniques. Comprehensive experimental evaluations indicate that the MTD-based mechanism significantly mitigates a range of poisoning attack types across multiple datasets with different federation topologies.
@@ I Introduction
This work proposes an MTD-based approach to dynamically modify the attack surface of DFL systems and improve their resilience and security. Therefore, the main contributions of this paper are: (1) Designing an MTD-based approach to mitigate poisoning attacks in DFL. This approach utilizes a dual-reputation score system based on model similarity and loss, adopts proactive and reactive modes, and incorporates dynamic topology and aggregation strategies. (2) Employing the DBSCAN clustering algorithm to adjust the triggering threshold for MTD dynamically. (3) Implementing a prototype of the proposed MTD method on a real DFL platform [11]. (4) Conducting extensive experiments to compare the proposed algorithm with state-of-the-art methods, including scenarios with fully connected, ring, and star topologies, as well as untargeted poisoning and backdoor attacks on MNIST, FashionMNIST, and CIFAR10 datasets. The results demonstrate that the proposed MTD-based algorithm not only effectively mitigates poisoning attacks, outperforming existing methods, but also improves model performance on non-IID data.
@@ II Background and Related Work
TABLE I provides an overview of the existing literature on applying MTD strategies in the context of Artificial Intelligence (AI) Security to address diverse cyber threats. In the traditional AI field, MTD has been employed to mitigate adversarial samples and poisoning attacks. [12] proposes an MTD approach to enhance Deep Neural Network (DNN) robustness by dynamically changing model parameters during inference. This makes it harder for attackers to craft compelling adversarial examples, improving resilience against white and black-box attacks. Morphence [13] is an MTD strategy generating diverse student models from a base model with random weight perturbations. Randomly selecting a student model for each inference reduces the adversarial attack success rate while maintaining accuracy. MT-MTD [14] presents an MTD framework protecting DNNs against backdoor attacks in edge computing. It divides training data, has an attacker train each dimension, and the defender randomly selects dimensions for consensus, increasing attack cost and maintaining accuracy. MTDeep [15] enhances neural network robustness using MTD by dynamically changing model architecture, making it harder for attackers to craft adversarial attacks.
However, there is limited research on utilizing MTD methodology to address the security challenges of FL systems. [16] enhances FL security with MTD by introducing an augmented dual-shuffle mechanism. This dynamically changes the attack surface by shuffling client selection and model aggregation, improving system security against various poisoning attacks. Voyager [10] employs a reactive MTD-based aggregation protocol to bolster the resilience of DFL against poisoning attacks. This is achieved by dynamically manipulating the connectivity of the network topology. However, its efficacy remains to be validated in non-IID scenarios.
@@ IV-B MTD Module
An MTD module can be characterized by a triplet $<C,M,T>$, where $C$ signifies the configuration space within which the MTD module is capable of moving, $M$ indicates the strategy employed for the action, and $T$ represents the timing function that determines the intervals at which actions are executed [19]. In essence, the design of an MTD module necessitates addressing three fundamental questions: (1) What to move? (2) How to move? and (3) When to move [20]?
What to Move? The threat model analysis delineated in Section III-B reveals that the attackers’ knowledge and capabilities are restricted to their local area. Thus, the attack surface for benign nodes mainly comes from the model shared by the malicious nodes and the aggregation functions. This paper posits that the configurations of neighboring connections in benign nodes and aggregation functions can be dynamically moved to mitigate the risks posed by poisoning attacks. By modifying the selection of neighboring nodes, benign nodes can effectively diminish the influence exerted by malicious nodes, thereby altering the network topology of the federation. Additionally, by varying the aggregation function, the complexity of executing a successful attack targeting a specific aggregation method can be significantly increased, thereby enhancing the system’s resilience against such threats.
@@
How to Move? ”How to move” refers to the principles and strategies used to alter system configuration elements to increase unpredictability and confuse attackers. This work develops two MTD strategies: the dynamic topology strategy and the dynamic aggregation algorithm strategy.
@@ IV-B2 Dynamic Aggregation Algorithm Strategy
When to Move? In the MTD module, ”when to move” pertains to selecting the optimal moment to execute the strategies, transitioning from the existing state to an alternative configuration. This work proposes two modes of MTD: proactive and reactive. In the proactive mode, the benign node modifies its connected neighbors and the aggregation algorithm in each round. Conversely, in the reactive mode, the benign node requires a triggering event before initiating an appropriate response.
In conventional reactive MTD systems, the trigger of a strategy is primarily contingent upon a pre-established threshold [10]. However, this pre-established threshold necessitates a prior knowledge, which is often difficult to achieve in practice. Therefore, this paper proposes a dynamic threshold determination algorithm to automate acquiring the threshold for triggering MTD strategies. As shown in Algorithm 3, this work utilizes the Density-Based Spatial Clustering of Applications with Noise (DBSCAN) [21] algorithm to dynamically find the model similarity threshold, loss threshold, and the MTD trigger.
@@ V-A Deployment Configuration
Dataset MNIST, FashionMNIST, CIFAR10
@@
  • MNIST [18] is a widely-used handwritten digits (0-9) dataset for image recognition, consisting of 60,000 training images and 10,000 testing images, each in gray-scale and sized 28x28 pixels. A Multilayer Perceptron (MLP) model is used, consisting of two fully connected hidden layers with 256x128. The MLP is trained for 3 epochs per federated round, optimized using Adam with a learning rate of $1\mathrm{e}{-3}$, and cross-entropy loss is applied.
  • FashionMNIST [22] is a dataset created as a more challenging alternative to MNIST for benchmarks. It contains 60,000 training and 10,000 testing 28x28 gray-scale images across 10 fashion categories (e.g., trousers and coats). A small CNN with two convolutional layers (32 and 64 filters) is employed, followed by two fully connected layers. The model is trained on FashionMNIST for 3 epochs per round using Adam optimizer with a learning rate of $1\mathrm{e}{-3}$.
  • CIFAR10 [23] is a benchmark dataset for ML, consisting of 50,000 training and 10,000 testing images (32x32 pixels) across 10 classes (e.g., airplane, automobile, bird). The colorful images and diverse classes make it more challenging than MNIST and FashionMNIST. The SimpleMobileNet [24] model is utilized. It starts with a convolutional layer (32 filters, 3x3 kernel), followed by max pooling. Then, five depthwise separable convolutional layers are used, followed by adaptive average pooling and a dense layer. The model is trained using Adam optimizer (learning rate $1\mathrm{e}{-3}$) for 5 epochs per round.
@@
Federation Setups. The experiments were performed using a federation with 10 nodes with 10 rounds. The initial network topology varied among fully connected ring and star structures. The network topology could change dynamically during training if the topology MTD was enabled.
@@ V-B Experiments
Non-IID $\alpha$ Agg. Fully Ring Star Fully Ring Star Fully Ring Star
@@
Figure 3: Average F1-Score Results for Label Flipping
@@
Given that the attack strategy employed in this experiment is a random weight attack, the dynamic aggregation algorithm fails to demonstrate its advantage. To summarize, the defense strategies based on MTD effectively mitigate model poisoning attacks within non-IID environments.
@@
Extreme Non-IID Scenario. To investigate the effectiveness of the algorithms proposed in this study within extreme non-IID environments, this experiment is established to evaluate the performance of the reference algorithms alongside the MTD algorithms. This assessment is conducted under various attack scenarios within a fully-connected network, utilizing a Dirichlet $\alpha$ parameter of 0.1, as depicted in Figure6.
Figure 6: Performance of the Algorithms against poisoning attack on three Datasets with Fully connected topology when Dirichlet $\alpha$=0.1
The results indicate a general decline in the performance of all models across the three datasets, attributable to the heightened levels of non-IID conditions. Furthermore, regarding label flipping and model poisoning attacks, the results reveal that the extent of non-IID does not significantly impact the effectiveness of the MTD algorithms. Specifically, the defenses employing the reactive dynamic topology strategy and those integrating both dynamic topology and dynamic aggregation functions outperform the other algorithms.
TABLE IV: F1 Score with Reactive Dynamic Topology MTD across Three Datasets, with a PNR Rate of 30%, Alpha Value of 0.5, and Fully Connected Topology
Non-IID $\alpha$ Attack MNIST FashionMNIST CIFAR10
0.5 Lable Flipping 0.896$\pm$0.024 0.784$\pm$0.099 0.494$\pm$0.042
Model Poisoning 0.879$\pm$0.021 0.799$\pm$ 0.039 0.511$\pm$0.069
Scalability. To investigate the scalability of the algorithm presented in this paper, an experimental setup is conducted utilizing 20 nodes with a fully connected topology. This scenario involved simulating an environment where 30% of the nodes are attacked. The performance of the reactive dynamic topology algorithm is evaluated under two specific attack types: label flipping and model poisoning. A Dirichlet $\alpha$ parameter of 0.5 is utilized for this assessment. The results are summarized in TABLE IV, indicating that the federated approach utilizing the MTD algorithm remains resilient against poisoning attacks in a large network. This observation underscores the adaptability and scalability of the proposed algorithms across networks of varying sizes.
@@ VI Discussion
Extensive experiments have demonstrated that the proposed MTD-based approach can effectively mitigate poisoning attacks in non-IID environments.
A. Reduced effectiveness in the backdoor. Nevertheless, despite this approach demonstrating superiority over existing state-of-the-art algorithms, there is potential for enhancement in addressing targeted attacks, such as backdoor attacks. To investigate the reasons behind the sub-optimal performance of the MTD-based approach in backdoor attacks, this work analyzes the distribution of reputation scores generated by benign nodes. It is conducted under conditions where the PNR is set at 10% and within a fully connected topology. The comparison is performed across three datasets, focusing on model poisoning and backdoor attack scenarios.
@@
B. Enhanced performance in non-IID settings. It is essential to highlight that the proposed solution enhances the model performance of DFL in non-IID environments, especially when using the dynamic topology strategy. The average F1-Score achieved by the proposed approach reaches 0.95 for the MNIST dataset, 0.83 for FashionMNIST, and 0.70 for CIFAR10 when using the reactive dynamic topology strategy, as illustrated in Figure3 and 4. These results indicate that the algorithm not only mitigates the impacts of poisoning attacks but also enhances model performance in non-IID settings. The effectiveness of the dynamic topology lies in its tendency to facilitate connections and aggregations among nodes that exhibit high similarity, diminishing the impact of non-IID data.
C. Trade-off. Regarding trade-off, the algorithm proposed in this paper increases the computational complexity compared to reference models, but the increased complexity is within polynomial complexity. However, the reference algorithms, including Krum, also need to calculate the distance of the models. Therefore, when evaluating the advantages in terms of security afforded by implementing MTD, this study posits that the additional computational complexity is justifiable. Regarding network overhead, the algorithms introduced herein do not necessitate any supplementary transmission of information, thus avoiding any additional burden on network overhead.
@@ VII Conclusion and Future Work
This work proposes an MTD-based approach to mitigate poisoning attacks in DFL, especially in non-IID environments. It introduces a dual-reputation score mechanism assessing model similarity and loss, with proactive and reactive MTD modes using dynamic topology and aggregation strategies. By leveraging the DBSCAN clustering algorithm, the approach dynamically adjusts thresholds, overcoming the limitations of fixed threshold defenses. The dynamic topology enhances node connections aligned with local models. Experiments on MNIST, FashionMNIST, and CIFAR10 datasets demonstrate the approach’s effectiveness against untargeted attacks and superior performance over state-of-the-art aggregation algorithms, offering a robust security framework for non-IID data.